# EmbAdvisor: Adaptive Cache Management for Sustainable LLM Serving

### v1

@@ Abstract.
As large language models (LLMs) become widely used, their environmental impact — especially carbon emission — has attracted more attention. Prior studies focus on compute-related carbon emissions. In this paper, we find that storage is another key contributor. LLM caching, which saves and reuses KV caches for repeated context, reduces operational carbon by avoiding redundant computation. However, this benefit comes at the cost of embodied carbon from high-capacity, high-speed SSDs. As LLMs scale, the embodied carbon of storage grows significantly.
To address this tradeoff, we present EmbAdvisor, a carbon-aware caching framework that selects the optimal cache size for LLM serving. EmbAdvisor profiles different LLM tasks and uses an Integer Linear Programming (ILP) solver to select cache sizes that meet SLOs while minimizing total carbon emissions. Overall, EmbAdvisor reduces the average carbon emissions of a Llama-3 70B model by 9.5 % under various carbon intensities over the non-adaptive cache scenario, and can save up to 31.2 % when the carbon intensity is low.
@@ 1. Introduction
Large language models (LLMs) become widely adopted across applications—from dialogue systems (OpenAI, 2023; DeepSeek, 2025, 2024) to healthcare (Peng et al., 2023) and developer tools (GitHub, 2024). Although LLMs demonstrate remarkable capabilities, they raise concerns about the environmental impact (Ding and Shi, 2024; Stojkovic et al., 2025; Faiz et al., 2024; Li et al., 2025; Shi et al., 2024; Nguyen et al., 2024; Li et al., 2023b, 2024; Samsi et al., 2023) — particularly in terms of carbon emissions measured in carbon dioxide equivalent (CO2e).
@@
Carbon emissions in LLM serving systems come from two main sources: operational carbon and embodied carbon. Operational carbon comes from the electricity consumed by computing. The amount of carbon emitted per kilowatt-hour (kWh) depends on the grid’s energy source mix, quantified by the carbon intensity (CI), measured in the unit of gCO2e/kWh (Maji et al., 2023). Renewable energy sources like solar and hydro have low carbon intensities, while non-renewable sources like coal and gas have high carbon intensities. Embodied carbon comes from the hardware manufacturing process (Wu et al., 2022; Gupta et al., 2022). Serving LLMs at scale is compute-intensive. It demands high power that leads to high operational carbon emissions, and at the same time, requires high-performance hardware like high-end GPUs and machine learning accelerators that have high embodied carbon. For example, serving a single LLM prompt can emit over 20 times more carbon than a conventional web search query (Griffiths, 2020; Wong, 2023).
@@
To reduce the carbon emissions of LLM serving, prior work has focused primarily on optimizing the compute side. For example, DynamoLLM (Stojkovic et al., 2025) improves energy efficiency to lower operational emissions; GreenLLM (Shi et al., 2024) and EcoServe (Li et al., 2025) reuse older GPUs to lower embodied emissions. However, these efforts largely overlook another major contributor to carbon impact in LLM serving systems—storage.
@@
While compute-related carbon emissions have been widely studied, the carbon emissions of storage remain an underexplored but critical component in sustainable LLM deployment. LLM tasks often involve long context to provide specific information for better generation quality, such as long chat history in multi-turn conversation and long documents in comprehension tasks. Processing these long inputs can require thousands of tokens before generation (Liu et al., 2024a). To address this, prior work (Gao et al., 2024; Liu et al., 2024a; Gao et al., 2025; Yao et al., 2025) has introduced context caching, which saves KV cache of contexts in fast storage and reuses them when needed again. This reduces redundant computation, lowers latency, and decreases resource usage—ultimately reducing operational carbon emissions.
@@
Despite the benefits, context caching requires high-capacity and high-speed storage such as SSDs, which come with embodied carbon emissions. Prior work (Tannu and Nair, 2023; Ke et al., 2024; Seagate, 2025; Bhagavathula et al., 2024; McAllister et al., 2024a, b; Dell Technologies, 2019) has shown that SSDs have high embodied carbon. They can contribute to over 75 % of datacenter servers’ embodied carbon emissions (Tannu and Nair, 2023; Dell Technologies, 2019). This means that caching does not come for free — it can significantly increase the overall carbon emission of LLM serving. However, reducing such emission is challenging as there is currently no standard method for carbon accounting in storage yet, let alone effective characterization or optimization strategies.
@@
In this work, we present EmbAdvisor, a carbon-aware caching framework to mitigate storage-related carbon emissions from caching in LLM serving systems. The tradeoffs between the increased embodied carbon and operational carbon savings make finding the optimal caching configuration challenging. To guide cache decisions, we characterize LLM serving caching behavior and make two key observations. First, higher loads, i.e., higher request rates, yield more latency improvements and carbon savings from caching, as latency from the prefill phase significantly reduces due to cache hits. Second, carbon emission savings depend on the carbon intensity of the grid. When a grid has a high carbon intensity, operational carbon from computation dominates; when a grid has a low carbon intensity, embodied carbon from cache storage (SSD in this study) dominates. These observations indicate that the optimal cache configuration depends on multiple factors — the time and region of the grid that affect carbon intensity, and LLM task type and load of request that affect the operational carbon savings from cache hits. EmbAdvisor adapts to these factors to optimize cache use for carbon efficiency.
To accommodate fluctuations in load and carbon intensity, we design an adaptive cache controller that dynamically selects the optimal cache configuration, which profiles the LLM system under various loads and cache sizes. At serving, EmbAdvisor uses this profile to select the optimal cache configuration based on the current load and real-time carbon intensity. This enables EmbAdvisor to balance the tradeoff between operational carbon savings (from reduced computation) and embodied carbon emissions (from storage).
LLM serving performance is typically measured by two metrics: Time To First Token (TTFT) and Time Per Output Token (TPOT). The Service Level Objectives (SLOs) specify acceptable thresholds for TTFT and TPOT, requiring the majority of requests (e.g., $\geq$ 90 %) to meet the timing constraints (Zhong et al., 2024; Stojkovic et al., 2025; Qin et al., 2024). Sometimes, the carbon-optimal choice may choose a smaller cache due to its lower embodied carbon. However, reducing cache size can hurt the hit rate, leading to worse performance. To ensure these SLOs are met while reducing carbon emissions from caching, EmbAdvisor uses an Integer Linear Programming (ILP) solver to find the optimal cache size while ensuring SLO attainment.
@@
In addition, we introduce a new cache replacement policy for EmbAdvisor, as conventional cache replacement policies such as Least Recently Used (LRU) do not optimize for carbon emissions. We find that longer reused context (i.e., more tokens upon a cache hit) yields more computation savings and thus reducing the operational carbon. On the other hand, the cache entry size varies, leading to different embodied carbon overheads. Therefore, we design a carbon-saving-aware (CSA) replacement policy that incorporates not only cache access frequency and recency, but also the overhead of embodied carbon and savings from reused context.
We implement EmbAdvisor on top of the open-source caching system, LMCache (LMCache Team, 2025b). EmbAdvisor leverages input SLO constraints and real-time carbon intensity to minimize total carbon emissions in serving LLMs while meeting the performance SLOs. We integrate two LLMs into EmbAdvisor: Llama-3 70B and 8B models (Meta, 2024), and evaluate on two tasks: multi-turn conversation using a ShareGPT dataset (Hugging Face, 2023) and document reading comprehension using the TriviaQA dataset (Joshi et al., 2017). Experiments are run on a system with 4 NVIDIA L40 GPUs and up to 16 TB of SSD storage.
@@
  • To the best of our knowledge, this is the first work that systematically studies the embodied carbon emission of storage due to caching in LLM serving systems.
  • We provide an adaptive solution that finds the carbon-optimal cache configuration for different carbon intensities and loads, while ensuring 90 % SLO attainment.
  • An evaluation indicates that EmbAdvisor lowers carbon emissions of a Llama-3 70B model by an average of 9.5 % under various carbon intensities, compared to a caching system without adaptivity. It can save up to 31.2 % when the carbon intensity is low. The carbon emission savings are prominent as LLMs are widely deployed in datacenters.
@@ 2. Background
In this section, we first introduce the LLM serving process. We then discuss cache mechanism for LLM serving for performance optimization. Finally, we discuss carbon accounting methods for operational and embodied emissions.
@@ 2.1. LLM Serving Process
In LLM serving systems, requests arrive continuously, so batching is key to maximizing throughput. Simply batches all incoming requests at once but does not allow new requests to join the batch mid-process leads to poor GPU utilization. To address this, recent work introduces continuous batching (Yu et al., 2022), which allows new requests to be inserted during the decode phase, significantly improving GPU efficiency.
@@ 2.2. Caching for LLM Serving
To further reduce the heavy computation in the prefill phase, recent work has introduced caching techniques (Gao et al., 2024; Liu et al., 2024a; Cheng et al., 2024; Yao et al., 2025; Kwon et al., 2023; Gao et al., 2025). These studies indicate that many requests share overlapping tokens. For example, in multi-turn conversations, later turns often include the entire chat history. In reading comprehension, multiple questions may ask about the same document. Caching enables reusing KV cache from previous requests, avoiding redundant computation for repeated context and trading off extra storage for better performance.
@@
Figure 1(a) illustrates a system that reuses KV cache of the context. Step ➊: upon a new request Prompt 1, the system processes it and saves its KV cache to storage. Step ➋: when the next request Prompt 2 arrives, the system looks up the cache for existing context. When hit, the system stitches the KV cache of the cached tokens and the new tokens, and processes them together. Figure 1(b) depicts the timeline with or without cache. Due to the intensive attention computation, KV cache lookup incurs a much lower overhead than processing. For example, the average TTFT in executing prompts in the ShareGPT dataset (Hugging Face, 2023) using the Llama-3 70B model on 4×4$\times$4 ×L40 GPUs is 1.7 seconds, but loading the KV cache of previous context only takes 0.03 second on average. Therefore, cache reduces TTFT significantly by eliminating redundant processing in the prefill phase. Even though caching does not reduce computation in the decode phase, it reduces the waiting time for the decode phase in continuous batching, improving the decode latency. Despite the performance gains, an effective caching system fundamentally relies on the storage backend to maintain a high volume of KV cache data that can be accessed at high speed. For example, caching 1000-token context for 1 million prompts of a Llama-3 70B model takes more than 300 TB (LMCache Team, 2025a).
@@ 2.3. Carbon Emission Accounting
Figure 2. Average carbon intensity (CI) and energy sources of four grids in 2024 (Electricity Maps, 2025).
@@
Storage 16 TB SSD 480 kgCO2e (Gupta et al., 2022)
LLM serving is highly compute-intensive and thus results in significant environmental impact. According to the carbon modeling tool from prior work (Gupta et al., 2022; Li et al., 2023a, b; Faiz et al., 2024; Shi et al., 2024; Nguyen et al., 2024; Li et al., 2024), the total carbon emissions of serving LLMs on a hardware platform are the sum of the operational ($C_{\rm o}$) and embodied ($C_{\rm e}$) carbon emissions. The embodied carbon emissions are derived from the execution time T and amortized over the system’s lifetime (LT). For example, the typical lifetime of hardware components is around 5 years (Gupta et al., 2022; Ostrouchov et al., 2020). In the rest of the paper, we follow the 5-year lifetime unless specified.
(1) $C=C_{\rm o}+\frac{\texttt{T}}{\texttt{LT}}C_{\rm e}.$
@@
The operational carbon emission is computed as the product of the energy $E$ consumed by running LLMs on the target hardware and the carbon intensity CI during the execution phase. Carbon intensity measures the amount of carbon dioxide equivalent emitted per unit of electricity generation (i.e., gCO2e/kWh). Lower carbon intensity value indicates a higher proportion of renewable energy used in electricity generation from the power grid. Figure 2 shows four regions – FR (France), ES (Spain), CISO (California in USA), and MISO (Midwestern and Southern states in USA) – that feature different sources of electricity generation.
(2) $C_{\rm o}=E\times\texttt{CI}.$
@@
The embodied carbon emission from hardware comes primarily from the semiconductor manufacturing process. In this work, we focus on the main hardware components that contribute to the majority of embodied carbon for LLM serving, including GPUs, CPUs, memory, and storage. Table 1 lists the hardware components and their embodied carbon ($C_{\rm e}$), which is modeled based on its processor chip area and memory capacity (Gupta et al., 2022). The total embodied carbon emission of the system running LLM serving is
(3) $C_{\rm e}=\sum_{{\rm comp}\in{\rm Platform}}C_{\rm e,comp}=C_{\rm e,GPU}+C_{\rm e,CPU}+C_{\rm e,Mem}+C_{\rm e,SSD}.$
@@
While prior studies have examined the embodied carbon of compute components like CPUs and GPUs, the impact of storage—particularly SSDs—for LLM serving remains underexplored, despite evidence that it contributes significantly to embodied emissions (Li et al., 2023a; Tannu and Nair, 2023; McAllister et al., 2024a; Bhagavathula et al., 2024; Dell Technologies, 2019; Seagate, 2025). In our evaluation platform, the embodied carbon from SSDs consume 76.6 % of the total embodied carbon of the server, similar to the over 75 % fraction reported by prior study (Tannu and Nair, 2023).
3. Performance and Carbon of LLM Caching

### v2

# Cache Your Prompt When It’s Green — Carbon-Aware Caching for Large Language Model Serving

@@ Abstract.
As large language models (LLMs) become widely used, their environmental impact, especially the carbon emission problem, has attracted more attention. Prior studies focus on compute-related carbon emissions. In this paper, we find that storage is another key contributor. LLM caching, which saves and reuses KV caches for repeated context, reduces operational carbon by avoiding redundant computation. However, this benefit comes at the cost of embodied carbon from high-capacity, high-speed SSDs. As LLMs scale, the embodied carbon of storage grows significantly. To address this tradeoff, we present GreenCache, a carbon-aware cache management framework that dynamically derives resource allocation plans for LLM serving. GreenCache analyzes the tradeoff between carbon emissions and SLO satisfaction, reconfiguring the resource over time to maintain this balance under dynamic workloads. Evaluations from real traces demonstrate that GreenCache achieves an average carbon reduction of 15.1 % when serving Llama-3 70B in the FR grid, with reductions reaching up to 25.3 %, while staying within latency constraints for ¿ 90 % of requests.
Sustainability, Large language model, context caching, carbon emissions
copyright: noneccs: Computer systems organizationccs: Computing methodologies Machine learningccs: Social and professional topics Sustainability
@@ 1. Introduction
Large language models (LLMs) become widely adopted across applications—from dialogue systems (OpenAI, 2023; DeepSeek, 2025; Google, 2024) to healthcare (Peng et al., 2023) and developer tools (GitHub, 2024). Although LLMs demonstrate remarkable capabilities, they raise concerns about the environmental impact (Ding and Shi, 2024; Stojkovic et al., 2025; Faiz et al., 2024; Li et al., 2025; Shi et al., 2024, 2025; Li et al., 2023b, 2024; Samsi et al., 2023), particularly in terms of carbon emissions measured in carbon dioxide equivalent (CO2e).
@@
Carbon emissions in LLM serving systems come from two main sources: operational carbon and embodied carbon. Operational carbon emissions come from the electricity consumed by computing. The amount of carbon emitted per kilowatt-hour (kWh) depends on the grid’s energy source mix, quantified by the carbon intensity (CI), measured in the unit of gCO2e/kWh (Maji et al., 2023). Renewable energy sources like solar and hydro have low carbon intensities, while non-renewable sources like coal and gas have high carbon intensities. Embodied carbon emissions come from the hardware manufacturing process (Wu et al., 2022; Gupta et al., 2022). Serving LLMs at scale is compute-intensive. It demands high power that will likely lead to high operational carbon emissions, and at the same time, requires high-performance hardware like high-end GPUs and machine learning accelerators that have high embodied carbon. For example, serving a single LLM prompt can emit over 20 times more carbon than a conventional web search query (Griffiths, 2020; Wong, 2023).
@@
To reduce the carbon emissions of LLM serving, prior work has focused primarily on optimizing the compute side (e.g., GPUs). For example, DynamoLLM (Stojkovic et al., 2025) improves energy efficiency to lower operational emissions; GreenLLM (Shi et al., 2024, 2025) and EcoServe (Li et al., 2025) reuse older GPUs to lower embodied emissions. However, these efforts largely overlook another major contributor to carbon impact in LLM serving systems: storage.
@@
While compute-related carbon emissions have been widely studied, the carbon emissions of storage remain an underexplored but critical component in sustainable LLM deployment. LLM tasks often involve a long context to provide specific information for better generation quality, such as long chat histories in multi-turn conversations and long documents in comprehension tasks. Processing these long inputs can require thousands of tokens before generation (Liu et al., 2024b). To address this, prior work (Gao et al., 2024; Liu et al., 2024b; Gao et al., 2025; Yao et al., 2025) has introduced context caching, which saves the KV cache of contexts in fast storage and reuses them when needed again. This reduces redundant computation, lowers latency, and decreases resource usage, ultimately reducing operational carbon emissions.
@@
Despite the benefits, context caching requires high-capacity and high-speed storage such as SSDs, which come with embodied carbon emissions. Prior work (Tannu and Nair, 2023; Ke et al., 2024; Seagate, 2025; Bhagavathula et al., 2024; McAllister et al., 2024a, b; Dell Technologies, 2019) has shown that SSDs contribute to over 75 % of datacenter servers’ total embodied carbon emissions (Tannu and Nair, 2023; Dell Technologies, 2019). This means that caching does not come for free, but can significantly increase the overall carbon emissions of LLM serving. In a cloud environment, storage is provisioned on demand and can scale. Therefore, the embodied carbon from service is counted by the storage size and its time of usage, like prior studies on embodied carbon of hardware (Gupta et al., 2022; Han et al., 2025). However, there is no effective characterization or optimization strategy that optimizes storage carbon for LLM serving.
@@
In this work, we present GreenCache, a carbon-aware caching framework to mitigate storage-related carbon emissions from caching in LLM serving systems. Balancing the tradeoff between increased embodied carbon and reduced operational carbon makes identifying the optimal caching configuration particularly challenging. To guide cache decisions, we characterize LLM serving caching behavior and make two key observations. First, higher loads amplify both latency reduction and carbon savings from caching, as cache hits substantially shorten the prefill phase. Second, carbon emission savings are strongly influenced by the carbon intensity (CI). When the carbon intensity becomes high, operational carbon from computation dominates; conversely, when it is low, embodied carbon from cache storage (SSD in this study) becomes the primary contributor. Both the load and CI fluctuate in a realistic environment, as load varies with user activity (Stojkovic et al., 2025), and CI changes with real-time energy sources (Maji et al., 2023). These insights reveal that the optimal cache configuration is jointly determined by these dynamic factors — the load, which governs operational carbon savings through latency reduction, and the carbon intensity, which affects the net carbon savings. Accordingly, GreenCache adapts to both load and CI conditions to continuously optimize cache configuration for optimal carbon efficiency.
GreenCache first profiles the performance and power under various workloads and cache sizes. In practice, GreenCache predicts both carbon intensity and workload based on historical values, enabling adaptive cache reconfiguration that anticipates future dynamics up to 24 hours in advance to preserve enough time to allow sufficient warm-up. However, finding the most carbon-efficient cache size is not sufficient. LLM serving applications are typically required to meet Service Level Objectives (SLOs) for both Time to First Token (TTFT) and Time to Produce Output Token (TPOT), with a target attainment rate (e.g., 90 %) to ensure service quality (Zhong et al., 2024; Stojkovic et al., 2025; Qin et al., 2024). To ensure SLO attainment while achieving carbon efficiency, we model this optimization as an Integer Linear Programming (ILP) problem. GreenCache integrates an ILP solver that takes the predicted carbon intensity and load, current TTFT and TPOT, and the SLOs as input, and determines the most carbon-efficient cache configuration that meets the SLO target, effectively balancing the operational carbon savings (from reduced computation) and embodied carbon emissions (from storage).
@@
In addition, we introduce a new cache replacement policy for GreenCache, as conventional cache replacement policies such as Least Recently Used (LRU) do not optimize for carbon emissions. We find that longer reused context (i.e., more tokens upon a cache hit) yields more computation savings, reducing the operational carbon. On the other hand, the cache entry size varies, leading to different embodied carbon overheads. Therefore, we design a new policy, Least Carbon Savings (LCS), that incorporates not only cache access frequency and recency, but also the embodied carbon costs and operational carbon savings from reused context.
[…]
@@
  • To the best of our knowledge, this is the first work that systematically studies the embodied carbon emissions of storage due to caching in LLM serving systems.
  • We design and implement a caching framework for LLM serving that dynamically reconfigures cache size in response to fluctuating CIs and workloads, while meeting the SLO attainment goal.
  • Our evaluation based on realistic request rate and CI traces demonstrates that GreenCache reduces average carbon emissions by 15.1 % in low-CI grids such as FR, and can still achieve up to a 6.91 % reduction in higher-CI grids like CISO when serving the Llama-3 70B model.
  • The source code of this work is available at https://greencache.persistentmemory.org.
@@ 2. Background
In this section, we first introduce the LLM serving process. We then discuss the cache mechanism for LLM serving for performance optimization. Finally, we discuss carbon accounting methods for operational and embodied emissions.
@@ 2.1. LLM Serving Process
In LLM serving systems, batching is a key technique for maximizing throughput (Kwon et al., 2023; Yu et al., 2022; Zhong et al., 2024). Naively batching all existing requests and disallowing batch enlargement mid-processing results in suboptimal throughput. To address this, recent work introduces continuous batching (Yu et al., 2022), which allows new requests to be inserted during the decode phase, significantly improving GPU efficiency.
@@ 2.2. Caching for LLM Serving
To further reduce the heavy computation in the prefill phase, recent work has introduced caching techniques (Gao et al., 2024; Liu et al., 2024b; Cheng et al., 2024; Yao et al., 2025; Kwon et al., 2023; Gao et al., 2025). These studies indicate that many requests share overlapping tokens. For example, in multi-turn conversations, later turns often include the entire chat history. In reading comprehension, multiple questions may ask about the same document. Caching enables reusing the KV cache from previous requests, avoiding redundant computation for repeated context and trading off extra storage for better performance.
@@
Figure 1(a) illustrates a system that reuses the KV cache of the context. Step ➊: upon a new request Prompt 1, the system processes it and saves its KV cache to storage. Step ➋: when the next request Prompt 2 arrives, the system looks up the cache for the existing context. When hit, the system stitches the KV cache of the cached tokens and the new tokens, and processes them together. Figure 1(b) depicts the timeline with or without cache. Due to the intensive attention computation, KV cache lookup incurs a much lower overhead than processing. For example, the average TTFT in executing prompts in the ShareGPT dataset (Hugging Face, 2023) using the Llama-3 70B model on $4\times$ L40 GPUs is 1.7 seconds, but loading the KV cache of previous context only takes 0.03 seconds on average. Therefore, cache reduces TTFT significantly by eliminating redundant processing in the prefill phase. Even though caching does not reduce computation in the decode phase, it reduces the waiting time for the decode phase in continuous batching, improving the decode latency. Despite the performance gains, an effective caching system fundamentally relies on the storage backend to maintain a high volume of KV cache data that can be accessed at high speed. For example, caching a 1000-token context for 1 million prompts of a Llama-3 70B model takes more than 300 TB (LMCache Team, 2025a).
@@ 2.3. Carbon Emission Accounting
Figure 2. (a) Average carbon intensity (CI) and energy sources of four grids in 2024 (Electricity Maps, 2025). (b) CI variation due to energy sources of the CISO grid on July 6, 2022 (Maji et al., 2023).
LLM serving is highly compute-intensive and results in significant environmental impact. According to the carbon modeling tool from prior work (Gupta et al., 2022; Li et al., 2023a, b; Faiz et al., 2024; Shi et al., 2024, 2025; Nguyen et al., 2024; Li et al., 2024), the total carbon emissions of serving LLMs on a hardware platform are the sum of the operational ($C_{\rm o}$) and embodied ($C_{\rm e}$) carbon emissions. As the cloud dynamically allocates resources, a user’s embodied carbon emissions are attributed to the duration for which the computing platform is allocated. Therefore, the embodied carbon emissions are proportional to the execution time T, amortized over the platform’s lifetime (LT). For example, the typical lifetime of hardware components is around 5 years (Gupta et al., 2022; Ostrouchov et al., 2020), which we follow in the rest of this paper unless specified. We calculate the total carbon emissions as
(1) $C=C_{\rm o}+\frac{\texttt{T}}{\texttt{LT}}C_{\rm e}.$
@@
Storage (Up to) 16 TB SSD (Up to) 480 kgCO2e (Gupta et al., 2022)
The operational carbon ($C_{\rm o}$) is computed as the product of the energy $E$ consumed by running LLMs on the target hardware and the carbon intensity (CI) during the execution
(2) $C_{\rm o}=E\times\texttt{CI},$
@@
where CI measures the amount of carbon dioxide equivalent emitted per unit of electricity generation (i.e., gCO2e/kWh). A lower carbon intensity value indicates a higher proportion of renewable energy used in electricity generation from the power grid. Figure 2a shows four grids: FR (France), FI (Finland), ES (Spain), and CISO (California, USA), which differ in their compositions of energy sources. Figure 2b shows the electricity sources vary in the CISO grid throughout a day.
@@
The embodied carbon ($C_{\rm e}$) from hardware comes primarily from the semiconductor manufacturing process. In this work, we focus on the main hardware components that contribute to the majority of embodied carbon for LLM serving, including GPUs, CPUs, memory, and storage. Table 1 lists the hardware components and their embodied carbon, which is modeled based on their processor chip area and memory capacity (Gupta et al., 2022). We calculate the total embodied carbon emissions of the system running LLM service as
(3) $C_{\rm e}=\hskip-11.38109pt\sum_{{\rm comp}\in{\rm Platform}}\hskip-11.38109ptC_{\rm e,comp}=C_{\rm e,GPU}+C_{\rm e,CPU}+C_{\rm e,Mem}+C_{\rm e,SSD}.$
@@
While prior studies have examined the embodied carbon of compute components like CPUs and GPUs, the impact of storage, particularly SSDs, for LLM serving remains underexplored, despite evidence that it contributes significantly to embodied emissions (Li et al., 2023a; Tannu and Nair, 2023; McAllister et al., 2024a; Bhagavathula et al., 2024; Dell Technologies, 2019; Seagate, 2025). In our platform, the embodied carbon from SSDs consumes 76.6 % of the total embodied carbon of the server, similar to the over 75 % fraction reported by a prior study (Tannu and Nair, 2023).
3. Performance and Carbon Emission of LLM Caching